OpenStudy (anonymous):

Find the equation of the circle with center at (3, 2) and through the point (5, 4).

OpenStudy (anonymous):

since the center is (3,2) so (x-3)^2+(x-2)^2=r^2 agree?

OpenStudy (anonymous):

sorry is (x-3)^2+(y-2)^2=r^2

OpenStudy (anonymous):

and (5,4) is on the circle , so put it into it then you'll get r=\[2\sqrt{2}\] so (x-3)^2+(y-2)^2=8

OpenStudy (anonymous):

Thank You

OpenStudy (anonymous):

welcome